OpenStudy (aaronandyson):

@ParthKohli 3 cos A = 4 sin A find cos A 3 - cot^2 A + cosec^2 A

OpenStudy (aaronandyson):

@Directrix

Directrix (directrix):

Is this question two problems or one? I get that these go together: 3 cos A = 4 sin A find cos A But, what is this: 3 - cot^2 A + cosec^2 A @AaronAndyson

OpenStudy (aaronandyson):

First 3 cos A = 4 sin A find cos A then 3 - cot^2 A + cosec^2 A

OpenStudy (aaronandyson):

@Directrix

OpenStudy (aaronandyson):

3 cos A = 4 sin A find 1. cos A 2.3 - cot^2 A + cosec^2 A

OpenStudy (aaronandyson):

@Directrix @satellite73 @ParthKohli

OpenStudy (aaronandyson):

@Directrix

OpenStudy (anonymous):

@AaronAndyson 1.cosA=? given that, 3cosA=4sinA => sinA/cosA=3/4 => tanA=3/4 =>tan^2 A=9/16 => sec^2 A -1 =9/16 => sec^2 A = 9/16 + 1 => sec^2 A = (9+16)/16 => sec^2 A = 25/16 => sec A = 5/4 [taking squ. root in both sides] =>1/cosA = 5/4 [since secA=1/cosA] => cosA = 4/5